OpenStudy (anonymous):

@zepdrix Find y" if y=8x sin x

OpenStudy (anonymous):

do you know product rule?

zepdrix (zepdrix):

Were you able to find the first derivative ok? :o

OpenStudy (anonymous):

I get confused when trigs are involved :( is it like 8 sin x + 8x cos x?

zepdrix (zepdrix):

\[\large y'=\color{royalblue}{(8x)'}\sin x+8x\color{royalblue}{(\sin x)'}\] Yah that looks right :)

OpenStudy (anonymous):

Awesome!

zepdrix (zepdrix):

\[\large y'=8 \sin x+8x \cos x\]Taking the derivative again, this first term shouldn't give you too much trouble, right? For the other term, you'll repeat this product rule process :O

OpenStudy (anonymous):

\[\int\limits (\int\limits (16 \text{Cos}[x]-8 x \text{Sin}[x]) \, dx) \, dx=8 x \text{Sin}[x] \]

zepdrix (zepdrix):

\[\large y''=\color{royalblue}{(8\sin x)'}+\color{royalblue}{(8x)'}\cos x+8x\color{royalblue}{(\cos x)'}\]

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

Okay so I got 8x sin x + 16 cos x! It should be -8x sin x so where does the negative come from?

zepdrix (zepdrix):

The derivative of that last term, cosine x should produce -sine x

OpenStudy (anonymous):

Got it! Okay just one last one I promise and you will be free of me!

zepdrix (zepdrix):

LOL no more XDDD

OpenStudy (anonymous):

LOL okay!